A boat goes downstream 72 miles in 3 hours, and travels upstream 60 miles in 6 hours. What is the rate of the river, and what is the boat in still water?
Answer by lwsshak3(11628)

A boat goes downstream 72 miles in 3 hours, and travels upstream 60 miles in 6 hours. What is the rate of the river, and what is the boat in still water?
***
let x=rate of boat in still water
let c= rate of the river
x+c=rate of boat downstream
x-c=rate of boat upstream
travel time*rate=distance
..
3(x+c)=72
6(x-c)=60
..
3x+3c=72
6x-6c=60
..
6x+6c=144
6x-6c=60
add
12x=204
x=17
..
subtract:
12c=84
c=7
rate of boat in still water=17 mph
rate of the river current=7 mph